OpenStudy (anonymous):

A grasshopper jumps at a 65.0 degree angle at 5.42m/s. At what time does it reach its maximum height?

OpenStudy (anonymous):

there is formula for finding time=2Visinangle/g..........2*5.42*sin65/9.8....=1.002sec..........for height=Vi[square]sin[squre]anlge/2g.....=1.21m

OpenStudy (whpalmer4):

@asmagul you should learn how to use the Equation button, it really isn't that difficult and makes a huge difference in readability. after you've observed what it does, you can type the appropriate commands directly, too.

OpenStudy (anonymous):

Time to reach maximum height is half of the time of flight\[vsin \theta / g\] v=5.42m/s θ=65